### v1

# Center-freeness of finite-step solvable groups arising from anabelian geometry

Abstract.
Anabelian geometry suggests that, for suitably geometric objects, their étale fundamental group determines the object up to isomorphism. From a group-theoretic viewpoint, this philosophy requires rigidity properties of the associated étale fundamental groups, which often follow from their center-freeness. In fact, some profinite groups arising from anabelian geometry are center-free. In the present paper, we investigate how such center-freeness behaves when passing to maximal $m$-step solvable quotients for any integer $m\geq 2$. In particular, we show that the maximal $m$-step solvable quotient of the geometric étale fundamental group of a hyperbolic curve over a field of characteristic 0 is center-free. Furthermore, we show that this implies the injectivity statement, i.e., the rigidity property, of the $m$-step solvable Grothendieck conjecture.
Key words and phrases:
étale fundamental group; anabelian geometry; hyperbolic curves; center-freeness; solvable quotients; Grothendieck conjecture
2020 Mathematics Subject Classification:
Primary 14H30; Secondary 14F35, 20E18
This work was supported by JSPS KAKENHI Grant Numbers 23KJ0881.
Introduction
Let $G$ be a profinite group. We define the derived series of $G$ by
$G^{[0]}\coloneq G,\qquad G^{[m]}\coloneq[G^{[m-1]},G^{[m-1]}]\qquad(m\geq 1).$
For each $m\in\mathbb{Z}_{\geq 0}$, we set $G^{m}\coloneq G/G^{[m]}$, and call it the maximal $m$-step solvable quotient of $G$. In the present paper, we consider the following property:
Property A.
Let $m\in\mathbb{Z}_{\geq 2}$. Then both $G$ and $G^{m}$ are center-free.
If $G$ is metabelian and center-free, then the natural projection $G\to G^{m}$ is an isomorphism for any $m\in\mathbb{Z}_{\geq 2}$, and hence A holds. In general, however, even if $G$ is center-free, $G^{m}$ need not be center-free. In fact, we can easily construct a counterexample as follows:
  • Let $D_{8}=\langle r,s\mid r^{4}=1,\ s^{2}=1,\ srs^{-1}=r^{-1}\rangle$ be the dihedral group of order 88, and define $\phi:D_{8}\to\operatorname{GL}_{2}(\mathbb{F}_{3})$ by
    $r\longmapsto\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\qquad s\longmapsto\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$.
    Then $G\coloneq(C_{3}\times C_{3})\rtimes_{\phi}D_{8}$ is center-free; however, $G^{2}\cong D_{8}$ is not center-free.
The following are known examples arising from anabelian geometry, which satisfy A:
  • Free pro-$\Sigma$ groups: Free pro-$\Sigma$ groups are center-free. Moreover, the maximal 22-step solvable quotients of free pro-$\Sigma$ groups are also center-free (see, for instance, [1, Section 4]). We can generalize this result from the case $m=2$ to any $m\in\mathbb{Z}_{\geq 2}$.
  • Absolute Galois groups: The absolute Galois groups of number fields and of $p$-adic local fields are center-free. Moreover, for any $m\in\mathbb{Z}_{\geq 2}$, their maximal $m$-step solvable quotients are also center-free (see [13, Proposition 1.1(ix) and Corollary 1.7]). This is closely related to the $m$-step solvable analogue of the Neukirch–Uchida theorem; see [13] for details.
In the present paper, we give a new example of such a profinite group arising from anabelian geometry. Let $k$ be a field of characteristic 0 with algebraic closure $\overline{k}$, and let $X$ be a smooth curve over $k$. Let $\Sigma$ denote a non-empty set of prime numbers. Write $\pi_{1}^{\mathrm{\acute{e}t}}(X)$ for the étale fundamental group of $X$. Moreover, for any $m\in\mathbb{Z}_{\geq 1}$, we define
$\Delta_{X}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{k}})^{\Sigma},\qquad\text{and}\qquad\Pi_{X}^{(m)}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(X)/\ker(\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{k}})\to\Delta_{X}^{m}).$
Then the following exact sequence holds:
$1\to\Delta_{X}^{m}\to\Pi_{X}^{(m)}\to G_{k}\to 1.$
If $X$ is hyperbolic, then $\Delta_{X}$ is center-free (see [14, Proposition 1.6]). For any profinite group $G$, we say that $G$ is ab-torsion-free if the abelianization of each open subgroup of $G$ is torsion-free. Additionally, we say that $G$ is ab-faithful if, for each open subgroup $H$ of $G$ and each open normal subgroup NN of $H$, the natural morphism
$H/N\rightarrow\operatorname{Aut}\bigl(N^{\mathrm{ab}}\bigr)$
induced by conjugation is injective. With the above notation, we have the following theorem:
Theorem A (Propositions 2.5 and 2.6).
Let $m\in\mathbb{Z}_{\geq 2}$. Then every ab-torsion-free, ab-faithful profinite group is center-free, and its maximal $m$-step solvable quotient is also center-free. In particular, if $X$ is hyperbolic, then both $\Delta_{X}$ and $\Delta_{X}^{m}$ are center-free.
The following is a direct corollary of this theorem:
Corollary A (Corollary 2.8).
Let $\mathcal{S}$ be a pro-$\Sigma$ surface groups of genus at least 22 and let $m\in\mathbb{Z}_{\geq 2}$. Then both $\mathcal{S}$ and $\mathcal{S}^{m}$ are center-free.
The same statement for free pro-$\Sigma$ groups of rank $\neq 1$ was proved in [15, Section 1.1]. However, the proof of [15, Proposition 1.1.1] contains an error and does not work as written. In Proposition 1.3, we correct the proof of [15, Proposition 1.1.1]; moreover, in Section 1 we give an explicit computation of the centralizer of a free generator. Throughout the present paper, for a profinite group $G$ and a closed subgroup $H\subset G$, we write $\mathrm{C}_{G}(H)$ for the centralizer of $H$ in $G$. We say that $G$ is slim if $\mathrm{C}_{G}(H)=1$ for every open subgroup $H$ of $G$. In particular, slimness implies that $G$ is center-free.
Theorem B (Theorems 1.5 and 1.6).
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$ with a free generating set $\mathcal{X}$. Let $m\in\mathbb{Z}_{\geq 2}$. Then, for any nonzero integer $n\in\mathbb{Z}$ and any $x\in\mathcal{X}$, one has
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle.$
In particular, $\mathcal{F}^{m}$ is slim if $r\neq 1$.
Next, we explain an application of A to the $m$-step solvable analogue of Grothendieck’s conjecture. The original Grothendieck conjecture was first proposed in a letter from Grothendieck to G. Faltings [5], and was proved by S. Mochizuki in [8]. Moreover, in [8, Theorem 18.1], S. Mochizuki proved the following “existence” statement for an $m$-step solvable analogue of the Grothendieck conjecture for hyperbolic curves over a sub-$p$-adic field (i.e., a field that embeds as a subfield of a finitely generated extension of $\mathbb{Q}_{p}$):
Assume $\Sigma=\{p\}$ and that $k$ is a sub-$p$-adic field. Let $m\in\mathbb{Z}_{\geq 2}$. Let $X_{1}$ and $X_{2}$ be smooth curves over $k$. Assume that at least one of $X_{1}$ and $X_{2}$ is hyperbolic. Then for any $G_{k}$-isomorphism
$\theta:\Pi_{X_{1}}^{(m+3)}\to\Pi_{X_{2}}^{(m+3)},$
there exists a $k$-isomorphism $\phi:X_{1}\to X_{2}$ such that the $G_{k}$-isomorphism $\Pi_{X_{1}}^{(m)}\to\Pi_{X_{2}}^{(m)}$ induced by $\phi$ (up to composition with an inner automorphism coming from $\Delta_{X_{2}}^{m}$) coincides with the isomorphism induced by $\theta$.
With a little additional argument, this theorem can be reformulated as the surjectivity of the following natural map:
We keep the notation and assumptions as above. Then the natural map
$\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)$ (0.1)
is surjective, where $\tilde{X}_{i}^{m}\to X_{i}$ is the maximal geometrically $m$-step solvable pro-$\Sigma$ Galois covering of $X_{i}$, and the right-hand set is the image of the natural map
$\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m+3)},\Pi_{X_{2}}^{(m+3)}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr).$
Then A induces the injectivity (i.e., “uniqueness”) statement as follows:
Corollary B (Theorem 2.9).
We keep the notation and assumptions as above. Then the natural map 0.1 is bijective.
In Section 1, we show B. In Section 2, we show A, A and B.
Notation and preliminaries in group theory
For any profinite group $G$, we define the derived series of $G$ by setting $G^{[0]}\coloneq G$ and, for any $m\in\mathbb{Z}_{\geq 1}$,
$G^{[m]}\coloneq[G^{[m-1]},G^{[m-1]}],$
where $\ast$$\ast$] denotes the closed subgroup topologically generated by commutators. Furthermore, for any $m\in\mathbb{Z}_{\geq 0}$, we set
$G^{m}\coloneq G/G^{[m]},$
and call it the maximal $m$-step solvable quotient of $G$. For simplicity, we write $G^{\mathrm{ab}}$ instead of $G^{1}$. Then we have the following lemma:
Lemma A (See [16, Lemma 1.1]).
Let $m,n\in\mathbb{Z}_{\geq 0}$. Let $G$ be a profinite group, and let $H$ be an open subgroup of $G^{m+n}$ containing $(G^{m+n})^{[m]}$. Let $\tilde{H}$ be the inverse image of $H$ in $G$ under the natural surjection $G\twoheadrightarrow G^{m+n}$. Then the natural surjection $\tilde{H}^{n}\twoheadrightarrow H^{n}$ is an isomorphism.
1. Centralizers in free $m$-step solvable groups
In this section, we compute explicitly the centralizer of a free generator of a free $m$-step solvable pro-$\Sigma$ group. A result of this form is stated in [15, Section 1.1]; however, the proof of [15, Proposition 1.1.1] contains an error and does not work as written. In Proposition 1.3 below, we provide a corrected argument. Throughout this section, let $\Sigma$ be a non-empty set of prime numbers.
Before turning to the proof, we record two ingredients.
$1.$ Pro-$\Sigma$ Fox calculus and the Blanchfield–Lyndon sequence
1.1.1.
Let us introduce the pro-$\Sigma$ Fox calculus and the Blanchfield–Lyndon sequence. For a pro-$\Sigma$ group $G$, we define its completed group ring by
[…]
where $H$ and $n$ run over all open normal subgroups of $G$ and all positive integers whose prime factors lie in $\Sigma$, respectively. In [2], R. H. Fox developed the (discrete) free differential calculus. Later, Y. Ihara [6] established a pro-$\Sigma$ analogue for a finitely generated free pro-$\Sigma$ group $\mathcal{F}$ with free generating set $X=\{x_{i}\}_{1\leq i\leq r}$. For any $i$, a continuous $\mathbb{Z}_{\Sigma}$-linear map
$\partial_{i}:\mathbb{Z}_{\Sigma}[[\mathcal{F}]]\to\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$
satisfying the following properties is called the free differential with respect to $x_{i}$:
  1. (i)
    $\partial_{i}(1)=0$, where 11 is the unit of $\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$;
  2. (ii)
    $\partial_{i}(x_{j})=\delta_{i,j}$;
  3. (iii)
    for any $\lambda,\tilde{\lambda}\in\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$, we have
[…]
    where $s$ is the augmentation morphism $\mathbb{Z}_{\Sigma}[[\mathcal{F}]]\to\mathbb{Z}_{\Sigma}$.
For any $i$, such a free differential is uniquely determined; see [6, Appendix]. Moreover, every $\lambda\in\mathbb{Z}_{\Sigma}[[\mathcal{F}]]$ admits an expansion
$\lambda=s(\lambda)\cdot 1+\sum_{i=1}^{r}\partial_{i}(\lambda)(x_{i}-1),$
and this expansion is unique (see [6, Theorem A-1]).
1.1.2.
Let $\mathcal{N}$ be a closed normal subgroup of $\mathcal{F}$. The conjugation action of $\mathcal{F}/\mathcal{N}$ on $\mathcal{N}^{\mathrm{ab}}$ extends continuously to an action of $\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]$. We regard $\mathcal{N}^{\mathrm{ab}}$ as a $\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]$-module by this action. Let $\pi:\mathbb{Z}_{\Sigma}[[\mathcal{F}]]\to\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]$ be the natural projection. For each $i$, define
$\tilde{\iota}:\mathcal{N}\rightarrow\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r};\qquad\tilde{\iota}(n)\coloneq\bigl(\pi\circ\partial_{i}(n)\bigr)_{1\leq i\leq r}.$
Since $\pi(n)=1$ for each $n\in\mathcal{N}$, we have $\tilde{\iota}(n_{1}n_{2})=\tilde{\iota}(n_{1})+\tilde{\iota}(n_{2})$. Therefore, $\tilde{\iota}$ is a homomorphism and factors through $\mathcal{N}^{\mathrm{ab}}$. Then we write $\iota$ for the induced morphism
$\mathcal{N}^{\mathrm{ab}}\rightarrow\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r}.$
Using the free differentials, Y. Ihara proved the profinite Blanchfield–Lyndon sequence:
Proposition 1.1 (The Blanchfield–Lyndon exact sequence, see [6, Theorem A-2]).
Let $\mathcal{F}$ be a free pro-$\Sigma$ group of finite rank $r$ with free generating set $X=\{x_{i}\}_{1\leq i\leq r}$, and let $\mathcal{N}$ be a closed normal subgroup of $\mathcal{F}$. Then there is an exact sequence
$0\rightarrow\mathcal{N}^{\mathrm{ab}}\xrightarrow{\iota}\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]^{\oplus r}\xrightarrow{f}\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]\xrightarrow{s}\mathbb{Z}_{\Sigma}\rightarrow 0$
of $\mathbb{Z}_{\Sigma}[[\mathcal{F}/\mathcal{N}]]$-modules, where $f$ is given by
$f\bigl((\lambda_{1},\cdots,\lambda_{r})\bigr)=\sum_{i=1}^{r}\lambda_{i}(\pi(x_{i})-1).$
The Blanchfield–Lyndon exact sequence admits a generalization to arbitrary profinite groups, known as the Complete Crowell exact sequence; see [10, Section 10.4] for details.
1.2. A first computation of a centralizer in a free $\Sigma pro$duct
1.2.1.
For a profinite group $G$ and an element $g\in G$, we define
$\mathrm{C}_{G}(g)\coloneq\{h\in G\mid hgh^{-1}=g\},$
and call it the centralizer of $g$ in $G$. (Note that this group is closed in $G$, and hence profinite.) A slightly different version of the following proposition first appeared in [11, Lemma 2.1.2], where it was used to prove the center-freeness of free discrete groups. We generalize it to our setting as follows:
Lemma 1.2.
[…]
$\overline{(\,\cdot\,)}:\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\rightarrow\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$
be the reduction morphism induced by $\mathbb{Z}/\ell^{\sigma}\mathbb{Z}\twoheadrightarrow\mathbb{Z}/\ell\mathbb{Z}$. If $E\in\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$ satisfies $\tilde{n}E=0$, then $\overline{E}=0$ in $\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$.
Proof.
We have
$\ker\bigl(\tilde{n}:\mathbb{Z}/\ell^{\sigma}\mathbb{Z}\to\mathbb{Z}/\ell^{\sigma}\mathbb{Z}\bigr)=\frac{\ell^{\sigma}}{\gcd(\tilde{n},\ell^{\sigma})}\cdot(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})=\ell^{\sigma-\operatorname{ord}_{\ell}(\tilde{n})}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z}).$
Applying this, we obtain
$E\in\ker\bigl(\tilde{n}:\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\to\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\bigr)=\ell^{\sigma-\operatorname{ord}_{\ell}(\tilde{n})}\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z}).$
Since $\sigma-\operatorname{ord}_{\ell}(\tilde{n})\geq 1$, we have
$\ell^{\sigma-\operatorname{ord}_{\ell}(\tilde{n})}\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\subseteq\ell\,\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z}).$
On the other hand, $\ell\,\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$ is exactly the kernel of the reduction morphism $\overline{(\cdot)}$. Thus $\overline{E}=0$. ∎
Proposition 1.3.
Let $\Omega=\mathcal{C}*P$ be the free $\Sigma pro$duct (see [12, Proposition 9.1.2]) of a procyclic pro-$\Sigma$ group $\mathcal{C}$, topologically generated by an element $x$, and a pro-$\Sigma$ group $P$. Let $m\in\mathbb{Z}_{\geq 2}$. Then, for any $n\in\mathbb{Z}$ such that $x^{n}\neq 1$ in $\mathcal{C}$, one has
$\mathrm{C}_{\Omega^{m}}(x^{n})\subset\langle x\rangle\cdot(\Omega^{m})^{[m-1]}$ (1.1)
as a subgroup of $\Omega^{m}$, where $\langle x\rangle$ denotes the closed subgroup of $\Omega^{m}$ generated by the image of $x$.
Proof.
Since $x^{-1}$ is also a topological generator of $\mathcal{C}$, we may assume that $n\geq 1$. To prove 1.1, it suffices to show that
$\rho\left(\mathrm{C}_{\Omega^{m}}(x^{n})\right)\subset\langle\rho(x)\rangle$ (1.2)
for any continuous surjection $\rho:\Omega^{m}\twoheadrightarrow G$ onto a finite group $G$ that factors through the natural projection $\Omega^{m}\twoheadrightarrow\Omega^{m-1}$. Since $\Omega=\mathcal{C}*P$, we have $\Omega^{\mathrm{ab}}\cong\mathcal{C}^{\mathrm{ab}}\times P^{\mathrm{ab}}\cong\mathcal{C}\times P^{\mathrm{ab}}$. In particular, the composition of the natural morphisms $\mathcal{C}\rightarrow\Omega\rightarrow\Omega^{m}$ is injective, and the family of surjections $\rho$ such that $\rho(x^{n})\neq 1$ is cofinal. Therefore, we may assume that $\rho(x^{n})\neq 1$.
To prove 1.2, it suffices to construct a profinite group $\tilde{G}$ and a factorization
${\Omega^{m}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho}$$\scriptstyle{\psi}$${\tilde{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${G}$
such that
$\phi\bigl(\mathrm{C}_{\tilde{G}}(\psi(x)^{n})\bigr)\subset\bigl\langle\phi\circ\psi(x)\bigr\rangle.$ (1.3)
Indeed,
$\rho\bigl(\mathrm{C}_{\Omega^{m}}(x^{n})\bigr)=(\phi\circ\psi)\bigl(\mathrm{C}_{\Omega^{m}}(x^{n})\bigr)\subset\phi\bigl(\mathrm{C}_{\tilde{G}}(\psi(x)^{n})\bigr)\subset\bigl\langle\phi\circ\psi(x)\bigr\rangle=\langle\rho(x)\rangle.$
Let $s$ be the order of $\rho(x)$ in $G$. Let $\ell\in\Sigma$. Let $\sigma\in\mathbb{Z}_{\geq 1}$ such that $\sigma>\operatorname{ord}_{\ell}(sn)$. Let
$G\hookrightarrow\operatorname{GL}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$
be the left regular permutation representation for some sufficiently large $u\in\mathbb{Z}_{\geq 1}$, and regard $G$ as a subgroup of $\operatorname{GL}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$ via this embedding. Define a group $\tilde{G}$ by
$\tilde{G}\coloneq\left\{\left(\begin{array}[]{cc}A&B\\ 0&C\end{array}\right)\in\operatorname{GL}_{2u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\,\middle|\,A\in G,\ B\in\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z}),\ C\in\langle\rho(x)\rangle\right\}.$
By construction, $\tilde{G}$ fits into the short exact sequence
$1\to(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})^{\oplus u^{2}}\to\tilde{G}\to G\times\langle\rho(x)\rangle\to 1.$
Since $\rho:\Omega^{m}\twoheadrightarrow G$ factors through $\Omega^{m}\twoheadrightarrow\Omega^{m-1}$, the group $G\times\langle\rho(x)\rangle$ is (m1)(m-1)-step solvable. Therefore, $\tilde{G}$ is an $m$-step solvable pro-$\Sigma$ group. The surjection $\Omega\to\Omega^{m}\xrightarrow{\rho}G$ extends to a morphism $\psi:\Omega\to\tilde{G}$, defined by
$x\mapsto\left(\begin{array}[]{cc}\rho(x)&\rho(x)\\ 0&\rho(x)\end{array}\right),\qquad p\mapsto\left(\begin{array}[]{cc}\rho(p)&0\\ 0&I_{u}\end{array}\right)\qquad\text{for all }p\in P.$
Hence the morphism $\psi:\Omega\to\tilde{G}$ factors through $\Omega\twoheadrightarrow\Omega^{m}$. We also denote by $\psi$ the induced morphism $\Omega^{m}\to\tilde{G}$. Then the morphisms $\psi$ and the natural projection $\phi:\tilde{G}\to G$ satisfy $\rho=\phi\circ\psi$.
Finally, we show the desired property 1.3. Let $y\in\mathrm{C}_{\tilde{G}}(\psi(x)^{n})$ and write
$y\coloneq\left(\begin{array}[]{cc}A&B\\ 0&C\end{array}\right)\in\mathrm{C}_{\tilde{G}}\left(\left(\begin{array}[]{cc}\rho(x)&\rho(x)\\ 0&\rho(x)\end{array}\right)^{n}\right)$
Then we have that
$y\cdot\psi(x)^{sn}=\left(\begin{array}[]{cc}A&B\\ 0&C\end{array}\right)\cdot\left(\begin{array}[]{cc}I_{u}&snI_{u}\\ 0&I_{u}\end{array}\right)\cdot\left(\begin{array}[]{cc}\rho(x)^{sn}&0\\ 0&\rho(x)^{sn}\end{array}\right)=\left(\begin{array}[]{cc}A&snA+B\\ 0&C\end{array}\right)$
and
$\psi(x)^{sn}\cdot y=\left(\begin{array}[]{cc}\rho(x)^{sn}&0\\ 0&\rho(x)^{sn}\end{array}\right)\cdot\left(\begin{array}[]{cc}I_{u}&snI_{u}\\ 0&I_{u}\end{array}\right)\cdot\left(\begin{array}[]{cc}A&B\\ 0&C\end{array}\right)=\left(\begin{array}[]{cc}A&B+snC\\ 0&C\end{array}\right)$
are the same. By comparing the top-right blocks, we obtain sn(AC)=0sn(A-C)=0 in $\mathrm{M}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})$. By Lemma 1.2, this implies $\overline{A}=\overline{C}$ in $\mathrm{M}_{u}(\mathbb{Z}/\ell\mathbb{Z})$. Since $G\hookrightarrow\operatorname{GL}_{u}(\mathbb{Z}/\ell^{\sigma}\mathbb{Z})\twoheadrightarrow\operatorname{GL}_{u}(\mathbb{Z}/\ell\mathbb{Z})$ is still injective, we conclude that $\phi(y)=A=C$ in $G$. Therefore, $\phi(y)\in\langle\rho(x)\rangle$, i.e., 1.3 holds. This completes the proof. ∎
1.3. Proof of the slimness of free $m$-step solvable groups
1.3.1.
Using the above ingredients, we compute explicitly the centralizer of a free generator of a (possibly infinitely generated) free $m$-step solvable pro-$\Sigma$ group, and obtain the slimness of such profinite groups.
Lemma 1.4.
Let $\mathcal{F}$ be a free pro-$\Sigma$ group of finite rank $r$ with free generating set $X$. Then for any non-zero integer $n\in\mathbb{Z}_{\neq 0}$ and any $x\in X$, the element $\overline{x}^{n}-1$ is a non-zero-divisor in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$, where $\overline{x}$ is the image of $x$ in $\mathcal{F}^{\mathrm{ab}}$.
Proof.
Denote by $\mathbb{Z}(\Sigma)_{\geq 1}$ the set of all positive integers whose prime factors lie in $\Sigma$. We may assume that $n\geq 1$ as $\overline{x}^{-n}-1=-\overline{x}^{-n}(\overline{x}^{\,n}-1)$ in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$. Let us show that if $y\in\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$ satisfies $(\overline{x}^{n}-1)y=0$, then $y=0$.
Since $\mathcal{F}^{\mathrm{ab}}$ is a free $\mathbb{Z}_{\Sigma}$-module of finite rank $r$, we may identify
$\mathcal{F}^{\mathrm{ab}}\,\cong\,H\times\mathbb{Z}_{\Sigma},$
where $H\cong\mathbb{Z}_{\Sigma}^{r-1}$ is the free abelian factor generated by the images of $X\setminus\{x\}$, and the factor $\mathbb{Z}_{\Sigma}$ corresponds to $\overline{x}$. Put
$A\coloneq\mathbb{Z}_{\Sigma}[[H]].$
For each $N\in\mathbb{Z}(\Sigma)_{\geq 1}$, let $C_{N}\coloneq\langle\overline{x}_{N}\mid(\overline{x}_{N})^{N}=1\rangle\cong\mathbb{Z}/N\mathbb{Z}$. Then, by the definition of the completed group algebra and the above decomposition, we have
[…]
Here, we may regard $\overline{x}$ as the projective limit of $(\overline{x}_{N})_{N}$.
Write $y_{N}$ for the image of $y$ in $A[C_{N}]$. Since $\{1,\overline{x}_{N},\cdots,(\overline{x}_{N})^{N-1}\}$ is an AA-basis of $A[C_{N}]$, there exists a unique $c_{i}^{(N)}\in A$ such that
$y_{N}=\sum_{i=0}^{N-1}c_{i}^{(N)}\overline{x}_{N}^{i}.$
The equation $(\overline{x}^{n}-1)y=0$ implies $((\overline{x}_{N})^{n}-1)y_{N}=0$ for all NN, and hence
$0=((\overline{x}_{N})^{n}-1)\left(\sum_{i=0}^{N-1}c_{i}^{(N)}\overline{x}_{N}^{i}\right)=\sum_{i=0}^{N-1}\bigl(c_{i-n}^{(N)}-c_{i}^{(N)}\bigr)\overline{x}_{N}^{i},$
where indices $\{i\}$ of $c_{i}^{(N)}$ are taken in $\mathbb{Z}/N\mathbb{Z}$. By AA-linear independence of $\{\overline{x}_{N}^{i}\}$, we obtain $c_{i-n}^{(N)}=c_{i}^{(N)}$ for all $i\in\mathbb{Z}/N\mathbb{Z}$. In other words, $c_{i}^{(N)}$ is constant on cosets of the subgroup $\langle n\rangle\subset\mathbb{Z}/N\mathbb{Z}$.
Let $n=n_{\Sigma}\cdot n_{\Sigma^{\prime}}$ be the unique decomposition such that $n_{\Sigma}\in\mathbb{Z}(\Sigma)_{\geq 1}$ and that $n_{\Sigma^{\prime}}$ is coprime to all primes in $\Sigma$. Fix $M\in\mathbb{Z}(\Sigma)_{\geq 1}$ such that $n_{\Sigma}\mid M$, and let $k\in\mathbb{Z}(\Sigma)_{\geq 1}$ be arbitrary. As $n_{\Sigma^{\prime}}$ and kMkM are coprime to each other, we have $\langle n\rangle=\langle n_{\Sigma}\rangle\subset\mathbb{Z}/kM\mathbb{Z}$. Hence we may apply the above result for NN to kMkM, which gives $c_{i-n_{\Sigma}}^{(kM)}=c_{i}^{(kM)}$ for all $i\in\mathbb{Z}/kM\mathbb{Z}$. Therefore, by $n_{\Sigma}\mid M$, we obtain
$c_{i}^{(kM)}=c_{i+M}^{(kM)}=\cdots=c_{i+(k-1)M}^{(kM)}$ (1.4)
for all $i\in\mathbb{Z}/kM\mathbb{Z}$. Let $\pi:A[C_{kM}]\to A[C_{M}],\ \overline{x}_{kM}\mapsto\overline{x}_{M}$, be the natural projection induced by $\mathbb{Z}/kM\mathbb{Z}\twoheadrightarrow\mathbb{Z}/M\mathbb{Z}$. By $\pi(\overline{x}_{kM})=\overline{x}_{M}$ and 1.4, we have
$y_{M}=\pi(y_{kM})=\sum_{i=0}^{kM-1}c_{i}^{(kM)}\,\pi\bigl(\overline{x}_{kM}^{i}\bigr)=\sum_{i=0}^{M-1}\Bigl(\sum_{j=0}^{k-1}c_{i+jM}^{(kM)}\Bigr)\overline{x}_{M}^{i}=\sum_{i=0}^{M-1}\bigl(k\cdot c_{i}^{(kM)}\bigr)\overline{x}_{M}^{i}.$
Comparing this with $y_{M}=\sum_{i=0}^{M-1}c_{i}^{(M)}\overline{x}_{M}^{i}$, we obtain
$c_{i}^{(M)}=k\cdot c_{i}^{(kM)}\in kA$
for all $i\in\mathbb{Z}/M\mathbb{Z}$. By running over all $k\in\mathbb{Z}(\Sigma)_{\geq 1}$ and using the fact $\bigcap_{k}kA=\{0\}$, we obtain $y_{M}=0$. Since the set $\{M\in\mathbb{Z}(\Sigma)_{\geq 1}\mid n_{\Sigma}\mid M\}$ is cofinal in $\mathbb{Z}(\Sigma)_{\geq 1}$, it follows that $y=0$. This completes the proof. ∎
Theorem 1.5.
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$ with free generating set $X$. Let $m\in\mathbb{Z}_{\geq 2}$. Then for any non-zero integer $n\in\mathbb{Z}_{\neq 0}$ and any $x\in X$, one has
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle.$
Proof.
If $r=1$, the assertion is clear. Hence we may assume $r\neq 1$. Fix $x\in X$. We divide the proof into three cases: the case $m=2$ with $r$ finite; the case of general $m$ with $r$ finite; and the general case.
First, we assume that $m=2$ and $r$ is finite. By Proposition 1.3 and the fact that $\langle x\rangle\subset\mathrm{C}_{\mathcal{F}^{2}}(x^{n})$, we obtain $\mathrm{C}_{\mathcal{F}^{2}}(x^{n})=\langle x\rangle\cdot(\mathrm{C}_{\mathcal{F}^{2}}(x^{n})\cap(\mathcal{F}^{2})^{[1]})$. Therefore, it suffices to show that
$\mathrm{C}_{\mathcal{F}^{2}}(x^{n})\cap(\mathcal{F}^{2})^{[1]}=1.$ (1.5)
Applying Proposition 1.1 to the case $\mathcal{N}=\mathcal{F}^{[1]}$, we obtain an injective $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$-linear morphism
$\iota:\ (\mathcal{F}^{2})^{[1]}\hookrightarrow\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]^{\oplus r}.$
Consider the conjugation action of $x^{n}$ on the abelian group $(\mathcal{F}^{2})^{[1]}$. By $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$-linearity of $\iota$, we obtain
$\mathrm{C}_{\mathcal{F}^{2}}(x^{n})\cap(\mathcal{F}^{2})^{[1]}$ $=$ $\{u\in(\mathcal{F}^{2})^{[1]}\mid x^{n}ux^{-n}=u\}$
$=$ $\ker\bigl((\overline{x}^{n}-1):(\mathcal{F}^{2})^{[1]}\to(\mathcal{F}^{2})^{[1]}\bigr)$
$\subset$ $\ker\bigl((\overline{x}^{n}-1):\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]^{\oplus r}\to\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]^{\oplus r}\bigr),$
where $\overline{x}$ is the image of $x$ in $\mathcal{F}^{\mathrm{ab}}$. By Lemma 1.4, the element $\overline{x}^{n}-1$ is a non-zero-divisor in $\mathbb{Z}_{\Sigma}[[\mathcal{F}^{\mathrm{ab}}]]$, and hence multiplication by $\overline{x}^{n}-1$ is injective. Therefore, the last kernel is trivial, and hence the equation 1.5 follows. This proves
$\mathrm{C}_{\mathcal{F}^{2}}(x^{n})=\langle x\rangle$
in the case where $r$ is finite.
Next, assume that $r$ is finite and proceed by induction on $m\in\mathbb{Z}_{\geq 2}$. The case of $m=2$ is already proved. Suppose that $m>2$ and that the assertion holds for m1m-1. As in the case $m=2$, by Proposition 1.3, it suffices to show that
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})\cap(\mathcal{F}^{m})^{[m-1]}=1.$ (1.6)
Let $g$ be an element of the left-hand side of 1.6. Let $H$ be an open normal subgroup of $\mathcal{F}^{m}$ containing $(\mathcal{F}^{m})^{[1]}$. Since
$\bigcap_{H}H^{[m-1]}=((\mathcal{F}^{m})^{[1]})^{[m-1]}=1,$
it suffices to show that $\rho_{H}(g)=1$ (i.e., $g\in H^{[m-1]}$) for each such $H$, where $\rho_{H}:H\twoheadrightarrow H^{m-1}$ is the natural surjection. The image of $x$ in the finite quotient $\mathcal{F}^{m}/H$ has finite order. Let NN denote this order. Since $g$ commutes with $x^{n}$, it also commutes with $x^{Nn}$, and therefore $\rho_{H}(g)$ commutes with $\rho_{H}(x^{Nn})$. By the Nielsen–Schreier theorem, the inverse image $\tilde{H}$ of $H$ in $\mathcal{F}$ is again a free pro-$\Sigma$ group, and we may choose a free generating set of $\tilde{H}$ containing $x^{N}$. By A, we have $H^{m-1}\cong\tilde{H}^{m-1}$. Applying the induction hypothesis for m1m-1 to $\tilde{H}^{m-1}$ and the basis element $x^{N}\in\tilde{H}^{m-1}$, we obtain
$\mathrm{C}_{H^{m-1}}\bigl((x^{N})^{n}\bigr)=\langle x^{N}\rangle.$
On the other hand, we have $g\in(\mathcal{F}^{m})^{[m-1]}\subset(\mathcal{F}^{m})^{[2]}\subset H^{[1]}$ and hence $\rho_{H}(g)\in(H^{m-1})^{[1]}$. Note that $\langle x^{N}\rangle$ maps injectively to $(H^{m-1})^{\mathrm{ab}}$, whereas $(H^{m-1})^{[1]}$ maps trivially. Therefore,
$\rho_{H}(g)\in\langle x^{N}\rangle\cap(H^{m-1})^{[1]}=1.$
This proves
$\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle$
in the case where $r$ is finite.
Finally, we consider the case $r=\infty$. Let $J$ be the directed set of finite subsets $X_{j}$ of $X$ such that $x\in X_{j}$. For each $j\in J$, let $\mathcal{F}_{j}$ be the finitely generated free pro-$\Sigma$ group on $X_{j}$ and let $\pi_{j}:\mathcal{F}\twoheadrightarrow\mathcal{F}_{j}$ be the continuous morphism sending generators in $X_{j}$ to themselves and generators in $X\setminus X_{j}$ to the identity element of $\mathcal{F}_{j}$. Additionally, let $\pi_{j}^{(m)}:\mathcal{F}^{m}\twoheadrightarrow\mathcal{F}_{j}^{m}$ be the natural projection induced from $\pi_{j}$. Then, by [12, Proposition 3.3.9], we have an isomorphism
[…]
Let $g\in\mathrm{C}_{\mathcal{F}^{m}}(x^{n})$. By the finite-rank case, we obtain $\pi_{j}^{(m)}(g)\in\langle\pi^{(m)}_{j}(x)\rangle$ for all $j$. Passing to the inverse limit, we conclude that $g\in\langle x\rangle$. Thus, $\mathrm{C}_{\mathcal{F}^{m}}(x^{n})=\langle x\rangle$ also holds when $r=\infty$. This completes the proof. ∎
For a profinite group $G$, we say that $G$ is slim if the centralizer $\mathrm{C}_{G}(H)$ of any open subgroup $H$ of $G$ is trivial. We note that slimness implies center-freeness.
Corollary 1.6.
Let $\mathcal{F}$ be a (possibly infinitely generated) free pro-$\Sigma$ group of rank $r$. Assume that $r\neq 1$. Then $\mathcal{F}^{m}$ is slim for all $m\in\mathbb{Z}_{\geq 2}$.
Proof.
Let $X$ be a free generating set of $\mathcal{F}$. Let $H$ be an open subgroup of $\mathcal{F}^{m}$, and take two distinct elements $x,x^{\prime}\in X$. Since $\mathcal{F}:H]<\infty$, there exist $n,n^{\prime}\geq 1$ such that $x^{n}\in H$ and $(x^{\prime})^{n^{\prime}}\in H$. Then Theorem 1.5 implies
$\mathrm{C}_{\mathcal{F}^{m}}(H)\subset\mathrm{C}_{\mathcal{F}^{m}}(x^{n})\cap\mathrm{C}_{\mathcal{F}^{m}}((x^{\prime})^{n^{\prime}})=\langle x\rangle\cap\langle x^{\prime}\rangle=1,$
where the last equality follows from the facts that $\langle x\rangle$ and $\langle x^{\prime}\rangle$ embed into the abelianization $\mathcal{F}^{\mathrm{ab}}$ and are distinct. This completes the proof. ∎
2. The $m$-step solvable Grothendieck conjecture
In this section, we show that the maximal $m$-step solvable quotients of the geometric étale fundamental groups of hyperbolic curves over a field of characteristic 0 are center-free (see Theorem 2.6). Moreover, we explain how to relate this result to anabelian geometry and the Grothendieck conjecture. Throughout this section, let $\Sigma$ be a non-empty set of prime numbers.
2.1. A strategy for $m$-step solvable center-freeness
2.1.1.
We record a strategy used in proofs of center-freeness for maximal $m$-step solvable quotients.
Definition 2.1 ([9, Definition 1.1]).
Let $G$ be a profinite group.
  1. (1)
    We say that $G$ is ab-torsion-free if, for each open subgroup $H$ of $G$, the abelianization $H^{\mathrm{ab}}$ is torsion-free.
  2. (2)
    We say that $G$ is ab-faithful if, for each open subgroup $H$ of $G$ and each open normal subgroup NN of $H$, the natural morphism
    $H/N\rightarrow\operatorname{Aut}\bigl(N^{\mathrm{ab}}\bigr)$
    induced by conjugation is injective.
  3. (3)
    Let $M$ be a (continuous) $G$-module. Then we say that the action $G\to\operatorname{Aut}(M)$ is fixed-point-free if $M^{G}=0$.
Remark $2.$
Let $G$ be a profinite group and let $m\in\mathbb{Z}_{\geq 2}$. For any open subgroup $P$ of $G$ such that $G^{[m-1]}\subset P$, let $\overline{P}\subset G^{m}$ be its image under $G\twoheadrightarrow G^{m}$. Then $P^{\mathrm{ab}}\xrightarrow{\sim}\overline{P}^{\mathrm{ab}}$ by A. In particular, the following hold:
  1. (1)
    Assume that $G$ is ab-torsion-free, and let NN be as in Definition 2.11. Then $\overline{N}^{\mathrm{ab}}$ is also torsion-free when $G^{[m-1]}\subset N$.
  2. (2)
    Assume that $G$ is ab-faithful, and let $H$ and NN be as in Definition 2.12. Assume that $G^{[m-1]}\subset N$. Then we have canonical isomorphisms
    $N^{\mathrm{ab}}\xrightarrow{\sim}\overline{N}^{\mathrm{ab}}\qquad\text{and}\qquad H/N\xrightarrow{\sim}\overline{H}/\overline{N}.$
    Hence the conjugation action of $\overline{H}/\overline{N}$ on $\overline{N}^{\mathrm{ab}}$ is also faithful.
Lemma 2.3.
Let $G$ be a profinite group and let $m\in\mathbb{Z}_{\geq 2}$. Assume the following two conditions:
  1. (a)
    The profinite group $G$ is ab-faithful;
  2. (b)
    The conjugation action of $G^{m}$ on $(G^{m})^{[m-1]}$ is fixed-point-free.
Then $G^{m}$ is center-free.
Proof.
Let NN be an open normal subgroup of $G$ containing $G^{[m-1]}$, and let $\overline{N}\subset G^{m}$ be its image under $G\twoheadrightarrow G^{m}$. By the definition of the center, we have
$Z(G^{m})\subset\ker\Bigl(G^{m}\to\operatorname{Aut}(\overline{N}^{\mathrm{ab}})\Bigr).$
The action of $G^{m}/\overline{N}$ on $\overline{N}^{\mathrm{ab}}$ is faithful by a (see Remark 2.22), and hence the right-hand kernel equals $\overline{N}$. By running over all such NN, we obtain
$Z(G^{m})\subset\bigcap\overline{N}=(G^{m})^{[m-1]}.$ (2.1)
On the other hand, for any $a\in G^{m}$, the condition $a\in Z(G^{m})$ is equivalent to the condition that $gag^{-1}=a$ for all $g\in G^{m}$, hence
$Z(G^{m})\cap(G^{m})^{[m-1]}=((G^{m})^{[m-1]})^{G^{m}}=1,$
where the last equality follows from b. Combining this with 2.1, we obtain $Z(G^{m})=1$. ∎
Lemma 2.4.
Let $G$ be an ab-torsion-free profinite group. Then the conjugation action of $G^{m}$ on $(G^{m})^{[m-1]}$ is fixed-point-free for all $m\in\mathbb{Z}_{\geq 2}$.
Proof.
Let $\mathcal{N}$ be the set of all open normal subgroups of $G$ containing $G^{[m-1]}$. Fix $N\in\mathcal{N}$. First, we claim that the natural morphism
$(N^{\mathrm{ab}})^{G/N}\to G^{\mathrm{ab}}$ (2.2)
is injective. Consider the corestriction and the transfer morphisms
$\mathrm{cor}_{N}:N^{\mathrm{ab}}\to G^{\mathrm{ab}}\qquad\mathrm{transfer}_{N}:G^{\mathrm{ab}}\to N^{\mathrm{ab}}.$
Let $G/N=\cup_{1\leq i\leq[G:N]}a_{i}N$ be a disjoint union of left-cosets with representatives $\{a_{i}\}_{i}$. Then, for any $n\in N$, we have $\mathrm{transfer}_{N}(\mathrm{cor}_{N}(n))=\sum(a_{i}^{-1}na_{i})$ on $N^{\mathrm{ab}}$, i.e.,
$\mathrm{transfer}_{N}\circ\mathrm{cor}_{N}=\sum_{a\in G/N}a\text{-conjugation}$
on $N^{\mathrm{ab}}$. In particular, the restricted morphism $(\mathrm{transfer}_{N}\circ\mathrm{cor}_{N})\mid_{(N^{\mathrm{ab}})^{G/N}}$ coincides with multiplication by $[G:N]$. Since $G$ is ab-torsion-free, $N^{\mathrm{ab}}$ is torsion-free. Hence $\mathrm{transfer}_{N}\circ\mathrm{cor}_{N}$ is injective on $(N^{\mathrm{ab}})^{G/N}$. Therefore, the restricted morphism $(\mathrm{cor}_{N})\mid_{(N^{\mathrm{ab}})^{G/N}}$, which is the morphism 2.2, is also injective. This completes the proof of the claim.
By taking the abelianization of the exact sequence $1\to N\to G\to G/N\to 1$, we have the exact sequence
$(N^{\mathrm{ab}})_{G/N}\to G^{\mathrm{ab}}\to(G/N)^{\mathrm{ab}}\to 1,$
where $(N^{\mathrm{ab}})_{G/N}$ stands for the module of $(G/N)$-coinvariants of $N^{\mathrm{ab}}$. By running over all $N\in\mathcal{N}$, we obtain that the sequence
[…]
is also exact (see [12, Proposition 2.2.4]). Since $G^{\mathrm{ab}}\xrightarrow{\sim}\left(G^{m-1}\right)^{\mathrm{ab}}$, the left-hand morphism of 2.3 is the zero map. The natural morphisms $(N^{\mathrm{ab}})^{G/N}\rightarrow(N^{\mathrm{ab}})_{G/N}\rightarrow G^{\mathrm{ab}}$ induce
[…]
This completes the proof. ∎
Proposition 2.5.
Let $G$ be an ab-torsion-free ab-faithful profinite group, and let $m\in\mathbb{Z}_{\geq 2}$. Then both $G$ and $G^{m}$ are center-free.
Proof.
The center-freeness of $G$ follows from ab-faithfulness. The center-freeness of $G^{m}$ follows from Lemmas 2.3 and 2.4. ∎
$2.$ Proof of the center-freeness of the maximal $m$-step solvable quotients of the geometric étale fundamental groups of hyperbolic curves
2.2.1.
For any connected, locally Noetherian scheme $\mathcal{X}$, we define
$\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X},\ast)$
to be the étale fundamental group of $\mathcal{X}$, where $\ast:\operatorname{Spec}(\Omega)\rightarrow\mathcal{X}$ denotes a geometric point of $\mathcal{X}$ and $\Omega$ denotes an algebraically closed field. The profinite group $\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X},\ast)$ depends on the choice of base point only up to inner automorphisms, and therefore we omit the choice of base point below. If, moreover, $\mathcal{X}$ is a geometrically connected $k$-scheme for some field $k$ of characteristic 0, then we set
$\Delta_{\mathcal{X}}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X}_{\overline{k}})^{\Sigma},\qquad\text{and}\qquad\Pi_{\mathcal{X}}\coloneq\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X},\ast)/\ker(\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X}_{\overline{k}})\to\pi_{1}^{\mathrm{\acute{e}t}}(\mathcal{X}_{\overline{k}})^{\Sigma}).$
In this case, we have the following exact sequence, called the homotopy exact sequence:
$1\to\Delta_{\mathcal{X}}\to\Pi_{\mathcal{X}}\to G_{k}\to 1.$
Additionally, we define
$\Pi_{\mathcal{X}}^{(m)}\coloneq\Pi_{\mathcal{X}}/\Delta_{\mathcal{X}}^{[m]}.$
Then the homotopy exact sequence naturally induces the following exact sequence:
$1\to\Delta_{\mathcal{X}}^{m}\to\Pi_{\mathcal{X}}^{(m)}\to G_{k}\to 1.$ (2.4)
2.2.2.
Next, we consider the case of smooth curves, where we always assume that smooth curves are geometrically connected. We say that a smooth curve $X$ over $k$ of type $(g,r)$ is hyperbolic if 22gr<02-2g-r<0, i.e., $(g,r)\notin\{(0,0),(0,1),(0,2),(1,0)\}$. When $r\geq 1$, we say that $X$ is affine. The basic fact about hyperbolicity is that
[…]
(see [14, Corollary 1.4]). If $X$ is affine, then $\Delta_{X}$ is a free pro-$\Sigma$ group; if $X$ is proper, then $\Delta_{X}$ is a pro-$\Sigma$ surface group.
Theorem 2.6.
Let $X$ be a hyperbolic curve over a field $k$ of characteristic 0. Let $m\in\mathbb{Z}_{\geq 2}$. Then $\Delta_{X}^{m}$ is center-free. If, moreover, $G_{k}$ is center-free, then $\Pi_{X}^{(m)}$ is also center-free.
Proof.
The last assertion follows from the first assertion and 2.4. Then we show the first assertion. We may assume $k=\overline{k}$. Since $X$ is hyperbolic, $\Delta_{X}$ is non-abelian. When $X$ is affine, $\Delta_{X}$ is a free pro-$\Sigma$ group and hence the assertion follows from Corollary 1.6. Therefore, we may assume that $X$ is proper, i.e., $\Delta_{X}$ is a pro-$\Sigma$ surface group of genus $g\geq 2$. By Proposition 2.5, it suffices to show that $\Delta_{X}$ is ab-torsion-free and ab-faithful. The known result [14, Corollary 1.2] implies that $\Delta_{X}^{\mathrm{ab}}$ is torsion-free and hence $\Delta_{X}$ is ab-torsion-free, since any open subgroup of $\Delta_{X}$ is also an étale fundamental group of a proper hyperbolic curve over $k$.
Next, we show that $\Delta_{X}$ is ab-faithful. Let $H$ be an open subgroup of $\Delta_{X}$ and NN an open normal subgroup of $H$. To prove ab-faithfulness, we may replace $\Delta_{X}$ by $H$ and assume that $H=\Delta_{X}$. Let $X_{N}\to X$ be the connected finite étale Galois covering corresponding to NN, with Galois group $\operatorname{Gal}(X_{N}/X)(\cong\Delta_{X}/N)$. Fix $\ell\in\Sigma$. Then the morphism
$\operatorname{Gal}(X_{N}/X)\rightarrow\operatorname{Aut}_{\mathbb{Q}_{\ell}}(H^{1}_{\mathrm{\acute{e}t}}(X_{N},\mathbb{Q}_{\ell}))$
is also injective. (This is a classical result; see, for instance, [7, Proposition 1.3].) On the other hand, the $\operatorname{Gal}(X_{N}/X)$-module $H^{1}_{\mathrm{\acute{e}t}}(X_{N},\mathbb{Q}_{\ell})$ is the $\mathbb{Q}_{\ell}$-linear dual of $N^{\mathrm{ab},\ell}\otimes_{\mathbb{Z}_{\ell}}\mathbb{Q}_{\ell}$, with the conjugation action of $\Delta_{X}/N$ (see [4, Exposé XI, Section 5]). Therefore, the composition of the natural morphisms
$\Delta_{X}/N\to\operatorname{Aut}(N^{\mathrm{ab}})\to\operatorname{Aut}_{\mathbb{Z}_{\ell}}(N^{\mathrm{ab},\ell})\to\operatorname{Aut}_{\mathbb{Q}_{\ell}}\bigl(N^{\mathrm{ab},\ell}\otimes_{\mathbb{Z}_{\ell}}\mathbb{Q}_{\ell}\bigr).$
is injective. This proves that $\Delta_{X}$ is ab-faithful. This completes the proof. ∎
Remark 2.7.
Let $k$ be a field finitely generated over $\mathbb{Q}$. Then $G_{k}$ is center-free [3, Proposition 19.2.6], and hence $\Pi_{X}^{(m)}$ is center-free for all $m\in\mathbb{Z}_{\geq 2}$ by Theorem 2.6. On the other hand, $\Pi_{X}^{(\mathrm{ab})}$ is also center-free, although $\Delta_{X}^{\mathrm{ab}}$ is not; see [16, Proposition 1.3] for details.
Corollary 2.8.
The maximal $m$-step solvable quotients of a pro-$\Sigma$ surface group of genus $g\geq 2$ are center-free for all $m\in\mathbb{Z}_{\geq 2}$.
Proof.
There exists a smooth proper curve over an algebraically closed field whose pro-$\Sigma$ étale fundamental group is isomorphic to the pro-$\Sigma$ surface group. Thus, the assertion follows from Theorem 2.6. ∎
2.3. Injectivity of the $m$-step solvable Grothendieck conjecture
2.3.1.
Let $i$ range over $\{1,2\}$. Let $m\in\mathbb{Z}_{\geq 1}$. Let $k$ be a field of characteristic 0 with algebraic closure $\overline{k}$, and let $X_{i}$ be a smooth curve over $k$. We write $\tilde{X}_{i}^{m}\to X_{i}$ for the maximal geometrically $m$-step solvable pro-$\Sigma$ Galois covering of $X_{i}$, which is a scheme over $\overline{k}$. We introduce the following non-standard notation for isomorphism sets:
  • We denote by
    $\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)$
    the set of all pairs
[…]
  • Let $n\in\mathbb{Z}_{\geq 0}$. We denote by
    $\operatorname{Isom}_{G_{k}}^{(m+n)}(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)})$
    the image of the natural map
    $\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m+n)},\Pi_{X_{2}}^{(m+n)}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr).$
    Moreover, we define
    $\operatorname{Isom}_{G_{k}}^{\operatorname{Out},(m+n)}(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}})\coloneq\operatorname{Isom}_{G_{k}}^{(m+n)}(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}})/\operatorname{Inn}(\Delta_{X_{2}}^{m}),$
    where $\operatorname{Inn}(\Delta_{X_{2}}^{m})$ denotes the subgroup of $\operatorname{Isom}_{G_{k}}(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}})$ consisting of inner automorphisms induced by conjugation by elements of $\Delta_{X_{2}}^{m}$.
With the above notation, S. Mochizuki proved the following result, which is called the $m$-step solvable Grothendieck conjecture for hyperbolic curves:
Theorem ([8, Theorem 18.1]).
Assume $\Sigma=\{p\}$. Let $i$ range over $\{1,2\}$. Let $m\in\mathbb{Z}_{\geq 2}$. Let $k$ be a sub-$p$-adic field (i.e., a field that embeds as a subfield of a finitely generated extension of $\mathbb{Q}_{p}$) with algebraic closure $\overline{k}$, and let $X_{i}$ be a smooth curve over $k$. Assume that at least one of $X_{1}$ and $X_{2}$ is hyperbolic. Then the natural map
$\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\rightarrow\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)$ (2.5)
is surjective.
The following theorem shows that the natural map is also injective:
Theorem 2.9.
We keep the notation and assumptions as in the above theorem. Then the natural map 2.5 is bijective.
Proof.
If $\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)$$\emptyset$, then the statement is tautological. Hence we may assume that $\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi^{(m)}_{X_{1}},\Pi^{(m)}_{X_{2}}\bigr)\neq\emptyset$. First, by Theorem 2.6, $\Delta_{X_{1}}^{m}$ is nontrivial and center-free if $X_{1}$ is hyperbolic. If $X_{1}$ is not hyperbolic, then $\Delta_{X_{1}}^{m}$ is abelian. Therefore, we can reconstruct whether $X_{1}$ is hyperbolic from $\Delta_{X_{1}}^{m}$. Hence we may assume that $X_{1}$ and $X_{2}$ are both hyperbolic. Next, by definition, there is an exact sequence:
$1\to\operatorname{Inn}(\Delta_{X_{2}}^{m})\to\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\to\operatorname{Isom}_{G_{k}}^{\operatorname{Out},(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\to 1.$
On the geometric side, we have an exact sequence:
$1\to\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\to\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\to\operatorname{Isom}_{k}(X_{1},X_{2})\to 1.$
Therefore, we obtain a commutative diagram with exact rows:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{\overline{k}/k}\bigl(\tilde{X}_{1}^{m}/X_{1},\tilde{X}_{2}^{m}/X_{2}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{k}(X_{1},X_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Inn}(\Delta_{X_{2}}^{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{G_{k}}^{(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Isom}_{G_{k}}^{\operatorname{Out},(m+3)}\bigl(\Pi_{X_{1}}^{(m)},\Pi_{X_{2}}^{(m)}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
By the definition of $\tilde{X}_{2}^{m}\to X_{2}$, we have a canonical identification
$\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\,\cong\,\Delta_{X_{2}}^{m}.$
By Theorem 2.6, $\Delta_{X_{2}}^{m}$ is center-free. Therefore,
$\ker(\operatorname{Aut}_{X_{2,\overline{k}}}\bigl(\tilde{X}_{2}^{m}\bigr)\twoheadrightarrow\operatorname{Inn}(\Delta_{X_{2}}^{m}))=\mathrm{C}_{\Pi_{X_{2}}^{(m)}}(\Delta_{X_{2}}^{m})$
is trivial. Hence the left-hand vertical arrow in the above commutative diagram is bijective. Moreover, the right-hand vertical arrow is surjective by [8, Theorem 18.1], and injective by [16, Lemma 4.9]. (Note that [16, Lemma 4.9] assumed that $k$ is a field finitely generated over $\mathbb{Q}$. However, the proof can be applied to the case where $k$ is a sub-$p$-adic field.) Thus, by the snake lemma, the middle vertical arrow is also bijective. This completes the proof. ∎
Remark 2.10.
In Theorem 2.9, we assumed that $\Sigma=\{p\}$, i.e., $\Sigma=1$. This assumption is not needed if we further assume $m\geq 3$. The author expects that, even for $m=2$, the same statement should hold for an arbitrary $\Sigma$. To prove this, we would need to check whether the proof of [8, Theorem 18.1] applies in this setting as well. At the time of writing, the author has not attempted this modification.
Acknowledgements
The author would like to express sincere gratitude to Prof. Akio Tamagawa for his invaluable assistance and insightful suggestions throughout this research.